A ball thrown horizontally from the top of a building 55m high strikes the ground at a point 35m for om the building. Find the (A) time to reach the ground (B) the initial speed of the ball, and (C) the velocity with the ball will strike the ground.
1
Expert's answer
2020-11-11T08:58:34-0500

Given:

"h=55\:\rm m;"

"l=35\:\rm m."

(A)

"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2\times 55}{9.8}}=3.35\: \rm s"

(B)

"v_0=\frac{l}{t}=\frac{35}{3.35}=10.4\:\rm m/s"

(C)

"v=\sqrt{v_0^2+2gh}=\sqrt{10.4^2+2\times 9.8\times 55}=34.5\:\rm m/s"